# Jinhao Zhang1  Yunquan Zhang2  Boyang Zhang3,4  Zeyu Liu5  Daning Cheng2

### v1

@@ Introduction
The experiments in this chapter will demonstrate the performance of the MoQE system. The experiments are divided into three parts. In CV tasks, the primary models are those based on residual architectures, specifically the ResNet and MobileNet series. In NLP tasks, our main focus is on current mainstream large language models, namely the Qwen and Llama series. Finally, we will analyze the relationship between the different system configurations and system performance, which we use as an ablation experiment.
CV Experiments
Experiments Configurations
The experiments configurations are listed as the following:
Model We use standard ResNet50, ResNet101, and MobileNetV2 as the experimental models. 8-bit integer (Int8) quantization is applied in this section.
Router Training Setting In training the router model, the same hyperparameter settings for the Adam optimizer were used across different backbone networks: the learning rate is set to $5.00\times 10^{-5}$ for all models, the learning rate schedule employs cosine annealing, the batch size was set to 48, gradient accumulation over 2 steps was applied, and weight decay was set to 0. Additional training details are provided in the Appendix.
Benchmarks and MoQE Setting. In our experiments, we use BRECQ (Li et al. 2021), DSConv (Nascimento, Prisacariu, and Fawcett 2020), N2UQ (Liu et al. 2021), and a QAT-trained model. The quantization expert models in MoQE consist of the above-mentioned benchmark models.
We present the performance of MoQE when the quantization expert models are not derived from the same full-precision model.
Experimental Resluts Analysis
Acc ResNet50 ResNet101 MobileNetV2
BRECQ 76.44% 77.60% 69.10%
QAT 77.09% 77.50% 71.10%
N2UQ 76.40% 78.30% 71.11%
DSConv 76.08% 77.90% 70.32%
MoQE(Our) 78.01% 78.91% 71.36%
FP16 80.21% 83.25% 71.80%
The experimental results are shown in Table  1. As can be observed from the experimental results in Table  1, the MoQE system demonstrates consistently superior performance across various quantization schemes on both ResNet and MobileNet series models. Overall, MoQE effectively mitigates the accuracy degradation typically induced by quantization while maintaining model efficiency.
NLP Experiments
Experiment Configurations
@@
Benchmarks and MoQE Setting In our experiments, we use SmoothQuant (Xiao et al. 2023), GPTQ (Frantar et al. 2022), Imatrix, and K-quantization. To evaluate the impact of the number of experts, we additionally include AWQ (Lin et al. 2024) as a benchmark in the configuration (Ablation) experiments. The quantization expert models in MoQE consist of the above-mentioned benchmark models.
@@
Experimental Results Analysis
@@
The experimental results are shown in Table  2. As shown in Table 2, the experimental results demonstrate the performance of the MoQE system across various large language models (Qwen0.6B, Qwen1.7B, Llama3B, and Qwen4B) and multiple NLP datasets (C4, WikiText-2, and OpenWebText). The evaluation metric is perplexity (lower is better), and MoQE is compared against several state-of-the-art quantization methods, including GPTQ, SmoothQuant, K-Quants, and imatrix, with FP16 serving as the full-precision baseline.
@@
Ablation Experiment
Int4 Experts
In this section, we will present the impact of various quantization level on the MoQE system. We use Int4 quantization models to build MoQE system. The experimental results are shown in Table  3.
@@
The experimental results indicate that the MoQE framework still exhibits significant advantages at the Int4 quantization level. Across all tested models and datasets, MoQE achieves the lowest perplexity, significantly outperforming each individual quantization method. Specifically, on the Qwen0.6B model, MoQE attains perplexities of 31.98 (C4) and 31.67 (WikiText-2), which are not only lower than those of GPTQ (33.12, 32.12), K-Quants (37.87, 36.53), imatrix (34.68, 34.15), and SmoothQuant (49.58, 47.21), but also closer to the FP16 baseline (17.18, 12.76). This indicates that MoQE can effectively mitigate the severe performance degradation caused by Int4 quantization in smaller models.
This advantage is equally pronounced in larger-scale models. On the Qwen1.7B, MoQE achieves perplexities of 21.72 (C4) and 20.44 (WikiText-2), surpassing all baseline methods, including the relatively better-performing imatrix (27.32, 27.89) and K-Quants (23.44, 25.32). Notably, on the Qwen4B model with even larger parameter sizes, MoQE’s advantage becomes more prominent, with perplexities of 17.09 (C4) and 13.89 (WikiText-2), which not only significantly outperform other quantization methods but also exhibit the smallest gap relative to the FP16 baseline (11.84, 7.95). This demonstrates the MoQE framework’s enhanced robustness and higher accuracy retention when handling large models under Int4 quantization.
@@
Table 4: MoQE system with Five Quantization Experts
As shown in Table 4, we extend the MoQE system by incorporating AWQ as an additional quantization expert, resulting in a total of five experts (GPTQ, SmoothQuant, K-Quants, imatrix, and AWQ). The performance of the individual AWQ expert and the aggregated MoQE system is analyzed below.
@@
More importantly, the MoQE system, which integrates routing over all five experts including AWQ, achieves the lowest perplexity across both models and datasets: 24.69 (C4) and 19.62 (WikiText-2) for Qwen0.6B, and 18.01 (C4) and 15.86 (WikiText-2) for Qwen1.7B. This demonstrates that the inclusion of AWQ as an expert further enhances the representational capacity of the ensemble, enabling the gating network to leverage complementary strengths across diverse quantization strategies. The superior performance of MoQE confirms that the dynamic expert selection mechanism effectively combines the best aspects of each method—particularly benefiting from AWQ’s strong baseline performance—thereby minimizing quantization error and outperforming even the best individual expert (AWQ) in all cases. This validates the scalability and robustness of the MoQE framework when extended to a larger set of heterogeneous quantization experts.
Conclusion

### v2

@@ 4 Experiments
The experiments are divided into three parts: NLP, CV, and Configurations experiment(Ablation experiments). In NLP tasks, our main focus is on current mainstream large language models, namely the Qwen and Llama series. In CV tasks, the primary models are those based on residual architectures, specifically the ResNet and MobileNet series. Finally, we will analyze how different system configurations affect performance, which we use as an ablation experiment.
4.1 NLP Experiments
4.1.1 Experiment Configurations
@@
Benchmarks and MoQE Setting In our experiments, we use SmoothQuant Xiao et al. (2023), GPTQ Frantar et al. (2022), Imatrix, and K-quantization. To evaluate the impact of the number of experts, we additionally include AWQ Lin et al. (2024) as a benchmark in the configuration (Ablation) experiments. The quantization expert models in MoQE consist of the above-mentioned benchmark models.
We present the performance of MoQE system when the quantization expert models are derived from the same full-precision model.
@@
4.1.2 Experimental Results Analysis
@@
As shown in Table 1, the experimental results demonstrate the performance of the MoQE system across various large language models (Qwen0.6B, Qwen1.7B, Llama3B, and Qwen4B) and multiple NLP datasets (C4, WikiText-2, and OpenWebText). The evaluation metric is perplexity (lower is better), and MoQE is compared against several state-of-the-art quantization methods, including GPTQ, SmoothQuant, K-Quants, and imatrix, with FP16 serving as the full-precision baseline.
@@
4.2 CV Experiments
4.2.1 Experiments Configurations
Model We use standard ResNet50, ResNet101, and MobileNetV2 as the experimental models. Int8 quantization is applied in this section.
Router Training Setting In training the router model, the same hyperparameter settings for the AdamW optimizer were used across different backbone networks: the learning rate is set to $5.00\times 10^{-5}$ for all models, the learning rate schedule employs cosine annealing, the batch size was set to 48, gradient accumulation over 2 steps was applied, and weight decay was set to 0. Additional training details are provided in the Appendix.
Benchmarks and MoQE Setting. In our experiments, we use BRECQ Li et al. (2021), DSConv Nascimento et al. (2020), N2UQ Liu et al. (2021), and a QAT-trained model. The quantization expert models in MoQE consist of the above-mentioned benchmark models.
4.2.2 Experimental Results Analysis
Acc ResNet50 ResNet101 MobileNetV2
BRECQ 76.44% 77.60% 69.10%
QAT 77.09% 77.50% 71.10%
N2UQ 76.40% 78.30% 71.11%
DSConv 76.08% 77.90% 70.32%
MoQE(Our) 78.01% 78.91% 71.36%
FP16 80.21% 83.25% 71.80%
@@ 4.2.2 Experimental Results Analysis
The experimental results are shown in Table  2. As can be observed from the experimental results in Table  2, the MoQE system demonstrates consistently superior performance across various quantization schemes on both ResNet and MobileNet series models. Overall, MoQE effectively mitigates the accuracy degradation typically induced by quantization while maintaining model efficiency.
4.3 Configuration Experiment
4.3.1 Int4 Experts
In this section, we will present the impact of various quantization level on the MoQE system. We use Int4 quantization models to build MoQE system. The results are shown in Table  3.
@@ 4.3.1 Int4 Experts
The results indicate that the MoQE still exhibits significant advantages at the Int4 quantization level. Across all tested models and datasets, MoQE achieves the lowest perplexity, significantly outperforming each individual quantization method. Specifically, on the Qwen0.6B model, MoQE attains perplexities of 31.98 (C4) and 31.67 (WikiText-2), which are not only lower than those of GPTQ, K-Quants, imatrix, and SmoothQuant, but also closer to the FP16 baseline. This indicates that MoQE can effectively mitigate the severe performance degradation caused by Int4 quantization.
This advantage is equally pronounced in larger-scale models. On the Qwen1.7B, MoQE achieves perplexities of 21.72 (C4) and 20.44 (WikiText-2), surpassing all baseline methods, including the relatively better-performing imatrix and K-Quants. Notably, on the Qwen4B model with even larger parameter sizes, MoQE’s advantage becomes more prominent, with perplexities of 17.09 (C4) and 13.89 (WikiText-2), which not only significantly outperform other quantization methods but also exhibit the smallest gap relative to the FP16 baseline. This demonstrates the MoQE framework’s enhanced robustness and higher accuracy retention when handling large models under Int4 quantization.
@@ 4.3.2 Number of Experts
Table 4: MoQE with five quantization experts
As shown in Table 4, we extend the MoQE system by incorporating AWQ as an additional quantization expert, resulting in a total of five experts.
@@
More importantly, the MoQE system, which integrates routing over all five experts including AWQ, achieves the lowest perplexity across both models and datasets: 24.69 (C4) and 19.62 (WikiText-2) for Qwen0.6B, and 18.01 (C4) and 15.86 (WikiText-2) for Qwen1.7B. This demonstrates that the inclusion of AWQ as an expert further enhances the representational capacity of the ensemble, enabling the router model to leverage complementary strengths across diverse quantization strategies. As shown in Fig 4, increasing the number of experts from 2 to 5 exhibits a trend of performance improvement for the MoQE system, indicating that greater expert diversity yields finer routing and consistently surpasses the best single quantization expert. The superior performance of MoQE confirms that the dynamic expert selection mechanism effectively combines the best aspects of each method—particularly benefiting from AWQ’s strong baseline performance—thereby minimizing quantization error and outperforming even the best individual expert (AWQ) in all cases. This validates the scalability and robustness of the MoQE framework when extended to a larger set of heterogeneous quantization experts.
4.4 End-to-End Inference Time and Memory Consumption Analysis
We measured the end-to-end inference latency of MoQE system to determine whether it would result in significant additional overhead. We primarily tested the Extra Inference Time (abbr. EIT), which includes the runtime of the router model and the time required to load the model into VRAM. We expect the unactivated quantized expert models to reside in main memory, while the activated experts are in GPU memory. Therefore, the additional inference time includes the time for loading experts and the time for the router. All inference-latency analyses were performed on a single NVIDIA V100S GPU.
Model Expert EIT EIT Ratio
Qwen 0.6B 2929.2ms 27.29ms 0.93%
Qwen 1.7B 5500ms 59.0ms 1.07%
LLaMA 3B 9364.8ms 222.68ms 2.38%
Qwen 4B 11606.4ms 157.4ms 1.36%
(a) Extra Inference Time ratio on NLP task.
Model Expert EIT EIT Ratio
ResNet-50 1.64ms 0.076ms 4.63%
ResNet-101 3.90ms 0.077ms 1.97%
MobileNetV2 0.925ms 0.074ms 8.0%
(b) Extra Inference Time ratio on CV task.
Table 5: Inference efficiency analysis of MoQE system.
As shown in Table 5 that The maximum EIT ratio remains within 8.0%, while the majority of the EIT ratio is below 3.0%. Such negligible latency confirms that MoQE increases accuracy without sacrificing the real-time responsiveness required for edge scenarios, fully preserving the efficiency advantages of the underlying quantized network. In MoQE system, the GPU keeps only one quantization expert and a lightweight router resident, the remaining experts are not co-resident and are loaded on demand. The peak GPU memory (VRAM) and host memory (CPU RAM) of MoQE system closely match those of a single quantization expert during inference. Detailed analyses of VRAM and CPU RAM are provided in the Appendix A.6.
5 Conclusion
@@
6 Reproducibility Statement
We provide an anonymized, downloadable code repository in https://github.com/paper-submission-goii/MoQE that includes complete scripts and instructions for constructing/loading quantization experts and for training/evaluation. The repository also contains environment files (exact dependency versions), data-preprocessing steps, and the hardware specifications used in our experiments, enabling reproduction of the main experimental results and ablations. All datasets are publicly available and used under their original licenses. Experiments code address is https://github.com/paper-submission-goii/MoQE.
Appendix A Appendix
A.1 Experiments Router Model Training Detail
A.1.1 CV Router Model Training Detail
In our CV router model training process, the primary objective is to end-to-end train a router model to learn how to intelligently and dynamically assign weights to a fixed set of experts based on the input data. To achieve fine-grained optimization of the router model, we employ the AdamW optimizer. We implement a staged dynamic learning rate strategy: during the initial 1 to 5 training epochs, the learning rate for the router model is set to 1e-5; from epochs 6 to 10, it is increased to 2e-5 to accelerate convergence; thereafter, it reverts to 1e-5 for fine-tuning. Additionally, an early stopping mechanism is implemented: if the validation accuracy does not improve for seven consecutive epochs, training will automatically terminate. This is complemented by a strategy to dynamically reduce the learning rate when accuracy plateaus, ensuring the model ultimately converges to its optimal performance.
A.1.2 NLP Router Model Training Detail
In the training of the router model, the model weights are initialized using a normal distribution, and the learning rate is set to $5\times 10^{-5}$. We employ the 8-bit AdamW optimizer, which has a lower memory footprint, combined with a cosine annealing learning rate scheduler featuring a 2000-step warm-up phase. The total loss function consists of two components: the standard cross-entropy loss and an adaptive load balancing loss. The weighting coefficient $\alpha$ for this auxiliary loss term is dynamically adjusted based on the degree of balance in expert activation during training (with a base value of 0.02); when imbalances in expert utilization occur, the penalty strength of this term increases accordingly, thereby encouraging the model to utilize all experts more uniformly. Regarding the training procedure, the model is trained for a total of 10 epochs. A curriculum learning strategy is adopted: during the first two epochs, the model is trained only on a subset of 70,000 samples; starting from the third epoch, it transitions to training on the full dataset, which is more challenging. We achieve an effective batch size of 48 by setting the per-device batch size to 8 and performing gradient accumulation over 6 steps, while applying gradient clipping with a maximum norm of 1.0 to ensure training stability.
A.2 The biases exhibited by different quantization models on CV tasks
This section supplements the observation that different Int8 quantization experts exhibit distinct biases in the data space for CV tasks. To validate this hypothesis, we constructed seven mutually exclusive sub-datasets from the ImageNet-1K dataset (each containing 5k–7k images with balanced label distributions). On this basis, we applied four mainstream Int8 quantization methods—BRECQ, QAT, N2UQ, and DSConv—to the ResNet-50 FP32 weights, ensuring that the data, hyper-parameters, and inference pipeline were completely identical, thereby excluding external variables from interfering with performance differences.
Figure 5: Top-1 accuracy of four Int8 quantization methods on seven ImageNet-1K sub-datasets.
The results are shown in the figure 5. By comparing the Top-1 accuracy of different quantization experts on each sub-dataset, we can see that for the same subset the performance of different quantization methods differs, and the relative performance ranking of each quantization model changes across different subsets; no single quantization method remains optimal on all subsets.For instance, on Subdataset3, N2UQ outperforms BRECQ by approximately 2.5% in Top-1 accuracy, whereas on Subdataset5, BRECQ surpasses N2UQ by 1.8%. The maximum gap across subsets reaches 4.7%, highlighting the strong data-dependent nature of quantization. This result empirically corroborates the core viewpoint proposed in Section 3.1 of the main text: that different quantization experts display bias distributions across the data, which provides the feasibility for dynamically assigning samples to the most suitable expert in the MoQE system.
A.3 routing probability assigned to each quantization expert
To better understand how the router model utilizes different quantization experts, we study the routing probabilities assigned during inference. For each sample, the router model will output a probability distribution about the quantization experts, which reflects the suitability of each quantization expert for the assigned sample. Despite differences between NLP and CV, both tasks demonstrate adaptive routing behaviors rather than static or random allocations. Figures 67 provide clear evidence of this phenomenon.
In NLP tasks (Figure 6), the router model displays strong sensitivity to input semantics. For certain datasets, the routing distribution is more concentrated, indicating that the router model consistently identifies one quantization expert as more reliable for specific linguistic patterns. In contrast, for other datasets, the probability mass is spread more evenly, suggesting that multiple quantization experts provide complementary strengths. These various probability distributions indicate that the router model can adaptively select the most suitable quantization expert based on the semantic context.
Figure 6: Routing probability distributions for NLP tasks.
In CV tasks (Figures 7), for deeper architectures such as ResNet101, routing probabilities reveal clearer specialization: inputs are more consistently assigned to a subset of experts, which aligns with the stronger bias of quantization methods on complex visual features. Conversely, for lightweight architectures such as MobileNetV2, the distributions are smoother and more balanced, implying that experts perform more comparably when the model has less representational complexity.
Figure 7: Routing probability distributions for CV tasks.
These results are consistent with the analysis in Appendix A.2, which demonstrated that different quantization models exhibit data-dependent biases. The routing probabilities show that the router model effectively captures these biases and adaptively allocates inputs to the most suitable expert. Overall, this probabilistic mechanism enables MoQE to integrate heterogeneous quantization methods and consistently achieve better performance than any single expert. These routing distributions further confirm that the proposed load-balancing design successfully prevents expert collapse during training, ensuring stable and diverse utilization of all experts.
A.4 The MoQE system with different numbers of quantization experts
This section provides additional analysis of MoQE under different numbers of quantization experts (2–5). As illustrated in Figures 8 , the results show a clear trend: increasing the number of experts from two to five brings improvements for both Qwen-0.6B and Qwen-1.7B, as the router model can utilize a richer library of quantization experts to make more accurate allocations. When adding AWQ as the fifth expert, MoQE system continues to improve. The MoQE system with five experts achieves the best results, surpassing the strongest single quantization expert, which confirms that MoQE system can effectively integrate the complementary advantages of different quantzation experts. Importantly, since each input only activates one quantization expert, the increase in the number of experts does not introduce significant inference overhead, which has been confirmed in the latency analysis. In summary, these data indicate that the performance of MoQE system steadily improves as the number of experts increases.
Figure 8: Impact of quantization expert size on MoQE system.
A.5 Convergence Observations and Analysis of the MoQE Router model
This section tracks the evolution of routing accuracy and Top-1 accuracy over the training process to analyze whether the router model stably assigns samples to the most suitable quantization expert, thereby validating the overall effectiveness of MoQE on CV tasks. In our setup, the expert weights remain frozen, and only the router model is fine-tuned, so as to avoid the confounding effect of “expert retraining” on the routing evaluation. We compute the corresponding metrics on the validation set throughout training and plot the curves to observe the convergence and robustness of the routing process.
Let the validation set be $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, and let $\{f_{j}\}_{j=1}^{M}$ denote $M$ frozen quantization experts. For a sample $(x_{i},y_{i})$, define the per-sample evaluation loss of expert $f_{j}$ as $\ell_{j}(x_{i},y_{i})$ (aligned with the task metric; lower is better). The optimal expert is
$j^{\star}(x_{i})=\arg\min_{j\in[M]}\,\ell_{j}(x_{i},y_{i}),\qquad[M]:=\{1,\ldots,M\}.$
Given a router model $r:\mathcal{X}\to[M]$, the Routing Accuracy is
$\mathrm{RA}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\!\left\{\,r(x_{i})=j^{\star}(x_{i})\,\right\}.$
Indicator function.
We denote by $\mathbf{1}\!\left\{\cdot\right\}$ the indicator function:
$\mathbf{1}\!\left\{A\right\}=\begin{cases}1,&\text{if the statement $A$ is true},\\ 0,&\text{otherwise}.\end{cases}$
Therefore, $\mathbf{1}\!\left\{\,r(x_{i})=j^{\star}(x_{i})\,\right\}=1$ if the router model picks the same expert as the optimal expert for sample $x_{i}$, and 0 otherwise.
To evaluate the router model’s learning process, we perform periodic evaluations on a fixed validation set throughout training. The system contains $N$ pre-quantized and frozen experts. We first annotate the optimal expert for each validation sample: for each sample $(x_{i},y_{i})$, we run a forward pass with all $N$ experts, compute the task-aligned evaluation loss, and record the expert with the smallest loss as that sample’s $j^{*}(x_{i})$. We then monitor two key metrics: (a) Routing Accuracy (RA), the proportion of samples for which the router’s choice $r(x_{i})$ matches the annotated optimal expert $j^{*}(x_{i})$; and (b) Top-1 accuracy, an end-to-end metric obtained by using the router selected expert $f_{r(x_{i})}$ to produce the prediction and comparing it with the ground-truth label $y_{i}$. We log both metrics at fixed intervals during training, thereby depicting the router’s approximation and convergence toward the “optimal expert” decision and its resulting impact on the system’s overall Top-1 accuracy.
Table 6 presents the system convergence upper bound measured by selecting the optimal expert per sample on the validation set. As shown, our router performs very well: across multiple models, the gap to this upper bound is small, and on MobileNetV2 the gap is only 0.15%.
Model System Convergence Upper Bound MoQE system $\bm{\Delta}$ to Upper Bound
ResNet-50 $\approx 78.44\%$ $78.01\%$ $0.43\%$
ResNet-101 $\approx 79.23\%$ $78.91\%$ $0.32\%$
MobileNetV2 $\approx 71.51\%$ $71.36\%$ $0.15\%$
Table 6: System convergence upper bound
As shown in Figure  9 (a), the curve exhibits a clear three-phase pattern: in the early stage of training, RA (routing accuracy) is about 25%; once training begins, it rapidly jumps to 60%. In the mid stage, it continues to rise steadily with small, controllable fluctuations (attributable to the trade-off between the load-balancing strategy). Finally, RA (routing accuracy) stably converges around  83%. Throughout this process, the quantization experts derived from the same base model are frozen and only the router is fine-tuned, which also demonstrates the effectiveness of the MoQE system. The improvement in RA (routing accuracy) can be directly interpreted as the router more frequently selecting the most suitable quantization expert for the task at hand, indicating that specialization matching to different sub-distributions is progressively taking shape.
(a) Routing Accuracy over training
(b) Top-1 accuracy over training (CV)
Figure 9: Routing convergence trends on CV tasks (a) Router Accuracy (RA) during training in the MoQE system. (b) Top-1 accuracy during training in the MoQE system
As shown in Figure 9 (b), The MoQE system’s Top-1 accuracy increases during training (from approximately 76% to 78%), with the rise of router accuracy in Figure 9 (a): the router more often assigns samples to the quantization expert that is better at the corresponding subdistribution, thus making the overall output distribution more aligned with the task objective and yielding a stable gain in Top-1 accuracy. This indicates that MoQE system’s gains indeed come from better routing rather than “stacking expert capacity”.
Figure 9 directly support the core proposition of the MoQE system: dynamic routing can enhance the accuracy of the quantization model. As training progresses, the probability that the router assigns samples to the most suitable quantization expert steadily rises to around 83%, and this is accompanied by a concurrent increase in the MoQE system’s Top-1 accuracy. The experiments in this section indicate that the Int8 expert models obtained through different quantization strategies based on the same base model show biases in the data distribution . The router learns from this and adaptively allocates the samples to the most suitable quantization experts.
A.6 Peak VRAM and CPU RAM USAGE ANALYSIS
This section focuses on runtime memory (the peak usage of CPU RAM and VRAM during inference or training). These metrics directly impact deployment feasibility and real-time performance (latency and throughput). Accordingly, we report and analyze peak VRAM and runtime CPU RAM to assess the practical deployment efficiency of MoQE system.
This subsection does not focus on offline data storage or disk space usage. Disk space usage has no direct impact on runtime memory or real-time performance and can be evaluated independently. Under MoQE system’s on-demand loading mechanism, the GPU keeps only a lightweight router and one Int8 expert resident , the remaining quantization experts are not resident in VRAM and are loaded only when an expert switch occurs. Meanwhile,disk storage has good scalability and large capacity—its capacity can even be measured in TB , which is significantly higher than the order of magnitude of VRAM and CPU RAM. We can increase disk capacity by upgrading local storage, mounting multiple disks, or using network-based storage solutions. Therefore, in common deployments, the size of offline weights does not constrain runtime memory or real-time performance , and this is not a significant obstacle for the MoQE system. Therefore, we consider the MoQE system’s greater disk space consumption to be an acceptable form of resource consumption.
As shown in Table 7,we report the storage size of model weights for different models under FP16/Int8/Int4. The router model’s weight size is very small: 16.7 MB (full precision) for the NLP router and 2.0 MB (full precision) for the CV router. Relative to the corresponding Int8 quantization experts, the router typically accounts for less than a few percent on large models for example, Qwen-1.7B: 1.70 GB ( 1% ), LLaMA-3B: 3.00 GB ( 0.6% ), Qwen-4B: 4.00 GB ( 0.4% ). On the CV tasks, ResNet-50: 25.6 MB ( 7.8% ) and ResNet-101: 44.5 MB ( 4.5% ). For very small backbones (e.g., MobileNetV2: 14.2 MB), the relative share increases, but the absolute router size still remains about 2 MB. Moreover, peak VRAM during inference is dominated by activations, so the router model’s extra cost in memory and latency is negligible. Combined with the MoQE system inference strategy (the GPU keeps only the lightweight router and the currently selected single Int8 quantization expert resident), MoQE system is on par with single expert Int8 in both storage and runtime memory, while achieving stable accuracy gains through better dynamic routing.
Category Model FP16 Int8 Int4
NLP Qwen-0.6B 1.20 GB 0.60 GB 0.30 GB
Qwen-1.7B 3.40 GB 1.70 GB 0.85 GB
LLaMA-3B 6.00 GB 3.00 GB 1.50 GB
Qwen-4B 8.00 GB 4.00 GB 2.00 GB
CV ResNet-50 51.2 MB 25.6 MB 12.8 MB
ResNet-101 89.0 MB 44.5 MB 22.3 MB
MobileNetV2 14.2 MB 3.4 MB 1.7 MB
Router NLP Router (full precision) 16.7 MB
CV Router (full precision) 2.0 MB
Table 7: Memory required to store the weights of different models.
Model VRAM (GB) CPU RAM (GB)
MoQE (Int8) Single(Int8) Single(FP16) MoQE Single (Int8/FP16)
Qwen-0.6B  $\approx\,4.32$  $\approx\,3.96$  $\approx\,4.62$  $\approx\,1.84$  $\approx\,0$ (slight)
Qwen-1.7B  $\approx\,7.86$  $\approx\,7.64$  $\approx\,9.43$  $\approx\,5.17$  $\approx\,0$ (slight)
LLaMA-3B  $\approx\,10.43$  $\approx\,10.17$  $\approx\,13.47$  $\approx\,9.43$  $\approx\,0$ (slight)
Qwen-4B  $\approx\,14.76$  $\approx\,14.30$  $\approx\,18.53$  $\approx\,12.23$  $\approx\,0$ (slight)
Table 8: Peak VRAM and CPU RAM during inference (NLP tasks).
Model VRAM(GB) CPU RAM
MoQE (Int8) Single(Int8) Single(FP16) MoQE Single (Int8/FP16)
ResNet-50  $\approx\,0.94$  $\approx\,0.90$  $\approx\,0.93$  $\approx\,76.8\,\text{MB}$  $\approx\,0$ (slight)
ResNet-101  $\approx\,1.35$  $\approx\,1.30$  $\approx\,1.34$  $\approx\,133.5\,\text{MB}$  $\approx\,0$ (slight)
MobileNetV2  $\approx\,0.38$  $\approx\,0.35$  $\approx\,0.36$  $\approx\,10.2\,\text{MB}$  $\approx\,0$ (slight)
Table 9: Peak VRAM and CPU RAM during inference (CV tasks).
Compared to a single quantization model, the incremental cost of MoQE system is concentrated in the CPU RAM and disk space . As shown in Table  8 and Table 9, the additional CPU RAM in NLP tasks is Qwen-0.6B: 1.84 GB, Qwen-1.7B: 5.17 GB, LLaMA-3B: 9.43 GB, Qwen-4B: 12.23 GB , in CV tasks, it is ResNet-50: 76.8 MB, ResNet-101: 133.5 MB, MobileNetV2: 10.2 MB. These values scale linearly with the size of the expert and occur only in CPU RAM and disk space. The aspect of disk space usage was previously analyzed.
Model Expert (ms) Router (ms) (I/O) (ms) I/O% Total% (Router+I/O)
Qwen-0.6B 2929.2 26.77 41.07 1.40% 2.32%
Qwen-1.7B 5500.0 57.60 121.45 2.21% 3.26%
LLaMA-3B 9364.8 220.05 212.78 2.27% 4.62%
Qwen-4B 11606.4 154.20 312.54 2.69% 4.02%
Table 10: Router model and expert loading overhead ratios.
The MoQE peak VRAM is effectively identical to the single Int8 baseline: for NLP, the difference is approximately 0.2 to 0.5GB, and for CV it is no more than 50MB, indicating no additional VRAM pressure. Relative to the FP16 baseline (Table 8), peak VRAM in NLP is substantially higher than with MoQE system as much as about 3.8 GB highlighting that, under the same batch size, peak VRAM is dominated by KV cache and intermediate activations rather than by weights. Under the same memory and latency budget, MoQE matches the single Int8 quantization model on runtime footprint and latency while surpassing it in Top-1 accuracy through higher routing accuracy that assigns each sample to the most suitable quantization expert. Relative to the FP16 baseline, MoQE system in can also reduce peak VRAM.
The MoQE system keeps one quantization expert resident in VRAM, with the remaining quantization experts resident in CPU RAM. The router model makes decisions on the GPU, if the selected quantization expert does not change, no weight switch is performed. Only when a switch is needed are the target quantization expert’s weights loaded from CPU RAM to VRAM (detailed data are provided in Table 10). Under this deployment strategy, the router model’s decision latency accounts for about 0.9%–2.4% of the quantization expert’s runtime, and the quantization expert loading latency (I/O) accounts for about 2%. Because a considerable portion of tasks do not need to switch quantization experts, the actual average cost of expert switching is lower. Even under a worst case switching scenario—where every request switches to a different quantization expert the end-to-end overhead remains within 5%, which we consider acceptable.
In summary , memory is not a bottleneck but an acceptable resource investment. MoQE system achieves no increase in VRAM and negligible speed impact, requiring only controlled CPU RAM (and corresponding disk space usage) to obtain stable accuracy gains.
A.7 Relevant Dataset, Models
To ensure the fairness and reproducibility of the experiment, a strict text purification process is carried out during the preprocessing stage: empty lines, HTML tags, URLs, and redundant lines containing repeated characters are removed, and all sequences are uniformly truncated to 2,048 tokens; the random seed is fixed throughout the process. Five mainstream quantization methods, namely AWQ, SmoothQuant, GPTQ, k-quants, and imatrix, are used in the experiment to quantize the four large language models, Qwen-0.6 B, Qwen-1.7 B, Llama-3 B, and Qwen-4 B, into quantization experts. The relevant weights have been fully open-sourced. The training and evaluation data are taken from three public datasets, C4, WikiText-2, and OpenWebText. All resources and processing scripts have been sourced from the official repository.
A.8 Use of LLMs
We used GPT-5 solely for language editing and refinement. All scientific ideas, experiments, and conclusions were conceived and verified by the authors, who take full responsibility for the content.